# Efficient and Transferable Agentic Knowledge Graph RAG via Reinforcement Learning

### v3

@@ B.2 Dataset Preprocessing
Following subgraph preprocessing practice RoG (Luo et al., 2024), we builds a question-specific near subgraph to shrink the search space and suppress spurious matches: starting from the linked anchor entities $e_{q}$ and gold answers $A_{q}$, it performs a breadth-first expansion over the KG $G$ up to the dataset’s maximum 2-hop radius $h$ The processed subgraph for each question is cached and used for KG retrieval server in KG-R1.
B.3 Freebase, Wikidata, and MultiTQ schema comparison
The following tables show the different schemas of Freebase, Wikidata, and MultiTQ.
@@ B.3 Freebase, Wikidata, and MultiTQ schema comparison
Freebase - P!nk - Gender - Ice Cube - United States of America - Nemanja Mikic - broadcast.content.artist - people.person.nationality - music.artist genre - location.location.containedby - basketball.basketball.player position
@@
WebQSP (4,737 test questions): Q: what does jamaican people speak A: [Jamaican English, Jamaican Creole English Language] CWQ (3,531 test questions): Q: Lou Seal is the mascot for the team that last won the World Series when? A: [2014 World Series] SimpleQA (21,687 test questions): Q: where is the madam satan located A: [United States of America] GrailQA (13,231 test questions): Q: which play is produced by the illusion? A: [The Illusion]
@@
Table 6: Temporal knowledge graph dataset used for KGQA evaluation. MultiTQ focuses on temporal reasoning with time-aware entities and relations.
Dataset Entities Relations
Temporal KG - Barack Obama (2009-2017) - World War II (1939-1945) - Steve Jobs (1955-2011) - Cold War (1947-1991) - Nelson Mandela (1994-1999) - president during - occurred during - CEO from to - started in - ended in
Questions
MultiTQ (9,000 test questions): Q: Who was the president of the United States when the iPhone was first launched? A: [George W. Bush] Q: Which major historical event ended the same year the European Union was established? A: [Cold War] Q: What technology company was founded during World War II? A: [IBM]
Appendix C Baselines
C.1 LLM-only Methods
@@
Setup. We evaluate Vanilla and CoT in a zero-shot setting without access to the KG retrieval server (i.e., no KG augmentation). All baselines run on Qwen-2.5-3B-IT, Qwen-2.5-7B-IT, and LLaMA-2-7B-Chat with temperature=0 and top_p=50..
Prompts. We use the following prompt templates for vanilla and CoT baselines.
@@
Prompte template for Vanilla setup Answer the given question directly and concisely based on your knowledge. Format your answer as: Answers: [”answer1”, ”answer2”, …]. For single answers, use: Answers: [”answer”]. Question: [Question] Answers:
@@
Prompte template for CoT setup Think through the question step by step, then provide the answer. IMPORTANT: Follow this exact format: 1. Start with ”Reasoning:” followed by your step-by-step thinking. 2. End with ”Answers:” followed by your final answer in brackets. 3. Do NOT put ”Answers:” before your reasoning. Format your answer as: Answers: [”answer1”, ”answer2”, …]. For single answers, use: Answers: [”answer”]. Question: [Question] Reasoning:
@@
Evaluation. Baseline outputs (Vanilla and CoT) differ in format from the KGQA gold answer set, so we employ an LLM-as-judge to align them. For each question, gpt-5-mini (OpenAI API) is given the question, the ground-truth answer set $A_{q}$, and parsed predited entities $\hat{A}_{q}$ from the base model’s output with a concise semantic-entity-matching prompt (Box below); it returns only a binary vector indicating, for each gold entity in order, whether the prediction refers to the same real-world entity at the same specificity (1 for exact/equivalent matches, e.g., “Apple Inc.” = “Apple”; 0 for overly general predictions, e.g., “Islam” vs [“Shia Islam”, “Sunni Islam”]). We report Pass@K (K=1,2,3,4), F1, precision, recall, and generation-token usage.
@@
Prompt template for LLM-as-Judge You are an evaluator performing semantic entity matching. Your task is to decide, for each gold entity (in order), whether the model’s prediction refers to the same real-world entity with the same level of specificity. Respond only with a binary vector in the format: $[0,1,0,1]$ Rules: - Exact matches or equivalent references → output 1 (e.g., ”Apple Inc.” = ”Apple”). - Too general or not specific enough when specificity is required → output 0 (e.g., ”Islam” vs [”Shia Islam”, ”Sunni Islam”]). Gold entities: [gold entity list] Predicted entities: [predicted entity list] Your Response:
C.2 LLM+KG baselinse
@@
RoG (Luo et al., 2024) Reasoning on Graphs (RoG) couples LLMs with KGs via a planning–retrieval–reasoning pipeline: the LLM first proposes KG-grounded relation paths as faithful plans, uses them to retrieve valid paths from the KG, and then performs stepwise reasoning to produce an answer. This yields interpretable multi-hop reasoning and reduces hallucinations by constraining reasoning to KG structure.
@@
ToG (Sun et al., 2024) Think-on-Graph (ToG) treats the LLM as an agent that iteratively explores the KG: it performs beam search over entities/relations, expands promising paths, and alternates retrieval with reasoning until a final path/answer is selected. Compared to prompt-only baselines, ToG improves deep, compositional reasoning by explicitly navigating the KG during multi-hop search.
@@
ReKnoS (Wang et al., 2025c) Reasoning with Knowledge Super-relations (ReKnoS) introduces super-relations that summarize and connect multiple relational paths, enabling both forward and backward reasoning while expanding the search space the LLM can traverse. By reasoning over these super-relations (rather than isolated edges), ReKnoS boosts retrieval success and multi-hop accuracy, especially on complex queries.
C.3 Cross-KG generliazation KG-specific Baselines
KG-Specific Baselines We selected one or two recent (2022–) state-of-the-art KG-specific methods, taking the authors’ reported test/dev scores as basline comparison. The below outlines baseline methods reported in Sec. 4.2 by KGQA benchmarks. The table 7 shows the summary.
[…]
Table 7: KG-specific Baselines For rows where Hits@1 is not reported, we show EM when available; where neither is reported, we show ”–”. Numbers are as reported in the cited papers/leaderboards.
[…]
GrailQA. SG-KBQA (Gao et al., 2025) is a schema-guided system that conditions a LLaMA-3.1-8B backbone on schema context and generates executable logical forms; official leaderboard reports overall Test EM/F1. ARG-KBQA (Tian et al., 2024) retrieves logical-form exemplars via an unsupervised ranker, then prompts GPT-3.5-0613 to generate and execute candidate logical forms.
QALD-10-en. COT-SPARQL (D’Abramo et al., 2025) applies chain-of-thought prompting with entity/relation hints to produce SPARQL; it reports both standard F1 and the Macro F1-QALD. DFSL-MQ (D’Abramo et al., 2024) performs dynamic few-shot retrieval with multi-query generation and beam selection (evaluated with CodeLlama-70B); it reports standard F1 on QALD-10-en.
T-REx. ATLAS (Izacard et al., 2023) is a retrieval-augmented seq2seq (FiD) reader paired with a dense retriever (Contriever); it reports Accuracyon T-Rex(treated as comparable to Hits@1).
MultiTQ (temporal KGQA). Prog-TQA (pro, 2024) uses in-context KoPL program generation with linking, execution, and a self-improvement loop; results are reported for LLaMA-13B/Vicuna-13B readers. ARI (Chen et al., 2024) (Abstract Reasoning Induction) separates planning (knowledge-agnostic) from execution (knowledge-based) with GPT-3.5-0613; it reports Accuracy (treated as comparable to Hits@1).
Notes: COT-SPARQL also reports Macro F1-QALD = 0.6387 on QALD-10 via GERBIL, while we list its standard F1 here for consistency across rows. ATLAS reports Accuracy on KILT hidden tests: zsRE = 80.8, T-REx = 85.1 (we do not map these to F1/Hits@1/EM). SG-KBQA numbers (EM=79.140, F1=84.403) are from the official GrailQA leaderboard (Test). ARG-KBQA is reported on Dev set (EM=59.6, F1=64.9). ARI reports Accuracy on MultiTQ (treated as comparable to Hits@1).
Appendix D KG-R1 details
D.1 Training algorithm
@@
Input: Dataset $D=\{(q,G,A_{q})\}$, base LLM $\pi_{0}$, horizon $H$, rollouts NN
@@
      Global: $R^{\text{global},(n)}=w_{\text{F1}}\cdot\text{F1}(\hat{A}_{q},A_{q})+w_{\text{ret}}\cdot v_{\text{ret}}$;
@@
D.2 Server Instruction Prompt
Server Instruction template consists of A,B,C. We use the same server instruction for both training and evaluation across experiments.
@@ D.2 Server Instruction Prompt
D.3 Reward weights
Our RL objective combines per-turn and episode-level signals with fixed weights. Per-turn rewards encourage well-structured interaction and effective retrieval: $r_{t}^{\mathrm{turn}}=w_{\mathrm{fmt}}\,s_{\mathrm{fmt}}+w_{\mathrm{kg}}\,s_{\mathrm{kg}}+w_{\mathrm{ans}}\,s_{\mathrm{ans}}$, where $s_{\mathrm{fmt}}$ scores output structure/format validity, $s_{\mathrm{kg}}$ rewards schema-valid, non-empty KG queries, and $s_{\mathrm{ans}}$ checks final-answer formatting/consistency. Episode-level reward emphasizes answer correctness and retrieval coverage: $R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}(\hat{A}_{q},A_{q})+w_{\mathrm{ret}}\cdot v_{\mathrm{ret}}$, with $v_{\mathrm{ret}}\in\{0,1\}$ indicating adequate retrieval coverage. Unless otherwise noted, we use the following Table 8 as defaults.
@@ D.4 RL Implementation & Hyperparameter
Batch Configuration and Gradient Accumulation. The 3B model uses a training batch size of $128$ with validation batch size of $256$, whereas the 7B model uses a training batch size of $256$ with validation batch size of $128$. During GRPO rollout, we collect $N=16$ trajectories per prompt for the 3B model and $N=8$ for the 7B model to balance exploration with memory constraints. The mini-batch size is $128$ for both models, with dynamic batch sizing enabled to utilize GPU memory efficiently.
@@
D.5 KG-R1 Retreival Server details
Setup. We implemented the KG-R1 retrieval server with FastAPI (Ramírez, 2018) and Uvicorn (Encode, 2018) to support (i) a schema-free 1-hop KG query API, (ii) high-throughput async batch execution, and (iii) robust validation and observability (regex action parsing, standardized <information> wrapping with auto-closure, and final-turn safeguards). After the KG-R1 agent generates a response, the parsed <kg-query> action is sent to the server, which performs exact string matching over the per-question knowledge graph to resolve entities and relations and returns the retrieved information. If the call is invalid—one of: KG_SERVER_ERROR: Invalid Action; KG_FORMAT_ERROR: Missing Required Fields; KG_FORMAT_ERROR: Wrong Argument Count; KG_SAMPLE_NOT_FOUND: Sample Missing; KG_ENTITY_NOT_FOUND: Entity Not in KG; KG_RELATION_NOT_FOUND: Invalid Relation; KG_NO_RESULTS: No Relations Found; KG_NO_RESULTS: No Entities Found—the server responds with a descriptive error message (see box below).
KG-R1 Server Error Examples KG_SERVER_ERROR: Invalid Action <error>Action ”get_entity_info” not available (use: get_head_relations, get_tail_relations, get_head_entities, get_tail_entities)</error>  KG_FORMAT_ERROR: Missing Required Fields <error>Missing required fields for get_tail_entities: relation_name</error>  KG_FORMAT_ERROR: Wrong Argument Count <error>get_tail_relations accepts only one entity argument</error>  KG_SAMPLE_NOT_FOUND: Sample Missing <error>Sample ”sample_12345” not found in KG</error>  KG_ENTITY_NOT_FOUND: Entity Not in KG <error>Entity ”Barack Obamaa” not found in KG</error>  KG_RELATION_NOT_FOUND: Invalid Relation <error>Relation ”location.capital” not found in KG</error>  KG_NO_RESULTS: No Relations Found <error>No tail relations found for entity ”Random_Entity_123” in knowledge graph</error>  KG_NO_RESULTS: No Entities Found <error>No tail entities found for relation ”film.director.film” with head ”Barack Obama” in knowledge graph</error>
@@ D.5 KG-R1 Retreival Server details
Appendix E Supplementary Experimental Results
Figure 6 shows Training dynamics of Qwen-2.5b-it are highly consistent across three random seeds, with smooth F1 improvements and minimal variance throughout 400 optimization steps. Both WebQSP (red) and CWQ (blue) curves show rapid early gains and stable convergence, indicating robust optimization and reproducible policy learning.
E.1 Reproducibility of KG-R1 Training
Figure 6: Training dynamics of Qwen 2.5b-it across 3 random seeds demonstrate reproducibility with steady F1 improvement and low variance. WebQSP (red) and CWQ (blue) metrics over 400 steps show stable convergence.
E.2 Latency and Throughput Analysis
@@
Batched Throughput We evaluate batched inference at batch size $64$ on a single NVIDIA H100 (Table 9). LLM-only baselines (no KG calls) achieve high throughput—$81.8$ (Vanilla) and $70.1$ (CoT) samples/s—driven by short generation (43.0/206.0 tokens per sample). KG-R1 incurs KG-server retrieval, reducing throughput but remaining practical for offline processing: the single-run setting reaches $3.7$ samples/s ($1205.9$ gen tokens/s) with $4.4$ KG calls per query, while the N=4N{=}4 runs setting trades throughput for more KG interaction ($17.5$ calls per query), yielding $2.0$ samples/s ($612.1$ gen tokens/s). Overall, KG-R1 sustains batch processing on one H100 while supporting KG-grounded reasoning.
@@
Configuration Samples Batch KG Calls Calls/Sample Total (s) Gen Tok./Sample Samples/s Gen Tok./s Vanilla Baseline 256 64 0 0.0 12.4 43.0 81.8 887.7 Vanilla CoT 256 64 0 0.0 14.1 206.0 70.1 3740.1 KG-R1 (single run) 256 64 1127 4.4 73.2 345.0 3.7 1205.9 KG-R1 (N=4N{=}4 runs) 256 64 4478 17.5 142.2 340.0 2.0 612.1
E.3 Ablation studies
All ablations in Table 3 evaluate KG-R1 with Qwen-2.5-3B-it trained on WebQSP and CWQ using a maximum turn budget of H=5H{=}5. We report the full ablation table in Table 3 (training curves in Figures 89).
Turn reward. We vary the turn-level reward by setting the weights to $w_{\text{fmt}}{=}0$, $w_{\text{kg}}{=}0$, and $w_{\text{ans}}{=}0$ (default: all $0.5$).
Turn-wise advantage. Instead of computing the turn-wise group advantage $A_{t}^{(n)}$ in Sec. 3.3, we compute a trajectory-wise group advantage

### v5

@@ C.2 Dataset Pre-processing
Following the subgraph pre-processing practice RoG [16], we build a question-specific near subgraph to shrink the search space and suppress spurious matches: starting from the linked anchor entities $e_{q}$ and gold answers $A_{q}$, it performs a breadth-first expansion over the KG $G$ up to the dataset’s maximum 2-hop radius $h$. The processed subgraph for each question is cached and used for the KG retrieval server in KG-R1.
C.3 Freebase and Wikidata Schema Comparison
The following tables show the different schemas of Freebase and Wikidata.
@@ C.3 Freebase and Wikidata Schema Comparison
Freebase - P!nk - Gender - Ice Cube - United States of America - Nemanja Mikic - broadcast.content.artist - people.person.nationality - music.artist_genre - location.location.containedby - basketball.basketball.player_position
@@
WebQSP (1,639 test questions): Q: What does jamaican people speak? A: [Jamaican English, Jamaican Creole English Language] CWQ (3,531 test questions): Q: Lou Seal is the mascot for the team that last won the World Series when? A: [2014 World Series] SimpleQA (21,687 test questions): Q: Where is the madam satan located? A: [United States of America] GrailQA (13,231 test questions): Q: Which play is produced by the illusion? A: [The Illusion]
@@
Appendix D Baselines
D.1 LLM-only Methods
@@
Setup. We evaluate IO, CoT, and SC in a zero-shot setting without access to the KG retrieval server (i.e., no KG augmentation). All baselines run on Qwen2.5-7B-it, Qwen2.5-72B-it, Qwen3-235B-A22B-Instruct-2507 with temperature=0 and top_p=50.
Prompts. We use the following prompt templates for IO and CoT baselines (the SC baseline performs majority voting over CoT results and hence uses the same prompt as CoT).
@@
Prompte template for IO Prompt setup Answer the given question directly and concisely based on your knowledge. Format your answer as: Answers: ["answer1", "answer2", …]. For single answers, use: Answers: ["answer"]. Question: [Question] Answers:
@@
Prompte template for CoT setup Think through the question step by step, then provide the answer. IMPORTANT: Follow this exact format: 1. Start with "Reasoning:" followed by your step-by-step thinking. 2. End with "Answers:" followed by your final answer in brackets. 3. Do NOT put "Answers:" before your reasoning. Format your answer as: Answers: ["answer1", "answer2", …]. For single answers, use: Answers: ["answer"]. Question: [Question] Reasoning:
@@
Evaluation. Baseline outputs (IO and CoT) differ in format from the KGQA gold answer set, so we employ an LLM-as-judge to align them. For each question, gpt-5-mini (OpenAI API) is given the question, the ground-truth answer set $A_{q}$, and parsed predicted entities $\hat{A}_{q}$ from the base model’s output with a concise semantic-entity-matching prompt (shown in the box below); it returns only a binary vector indicating, for each gold entity in order, whether the prediction refers to the same real-world entity at the same specificity (1 for exact/equivalent matches, e.g., “Apple Inc.” = “Apple”; 0 for overly general predictions, e.g., “Islam” vs [“Shia Islam”, “Sunni Islam”]). We report Pass@K (K=1,2,3,4), F1, precision, recall, and generation-token usage.
@@
Prompt template for LLM-as-Judge You are an evaluator performing semantic entity matching. Your task is to decide, for each gold entity (in order), whether the model’s prediction refers to the same real-world entity with the same level of specificity. Respond only with a binary vector in the format: $[0,1,0,1]$ Rules: - Exact matches or equivalent references → output 1 (e.g., "Apple Inc." = "Apple"). - Too general or not specific enough when specificity is required → output 0 (e.g., "Islam" vs ["Shia Islam", "Sunni Islam"]). Gold entities: [gold entity list] Predicted entities: [predicted entity list] Your Response:
D.2 KG-augmented Baselines
@@
RoG [16] Reasoning on Graphs (RoG) couples LLMs with KGs via a planning–retrieval–reasoning pipeline: the LLM first proposes KG-grounded relation paths as faithful plans, uses them to retrieve valid paths from the KG, and then performs stepwise reasoning to produce an answer. This yields interpretable multi-hop reasoning and reduces hallucinations by constraining reasoning to the KG structure.
@@
ToG [23] Think-on-Graph (ToG) treats the LLM as an agent that iteratively explores the KG: it performs beam search over entities/relations, expands promising paths, and alternates retrieval with reasoning until a final path/answer is selected. Compared to prompt-only baselines, ToG improves deep, compositional reasoning by explicitly navigating the KG during multi-hop search.
PoG [4] Plan-on-Graph (PoG) extends agentic KG reasoning with adaptive planning and self-correction. It decomposes a question into sub-objectives and iteratively performs path exploration, memory updating, and reflection to dynamically adjust reasoning trajectories. Through guidance, memory, and reflection mechanisms, PoG enables flexible exploration breadth and backtracking over erroneous paths, improving both reasoning accuracy and efficiency compared to fixed exploration strategies.
@@
ReKnoS [28] Reasoning with Knowledge Super-relations (ReKnoS) introduces super-relations that summarize and connect multiple relational paths, enabling both forward and backward reasoning while expanding the search space that the LLM can traverse. By reasoning over these super-relations (rather than isolated edges), ReKnoS boosts retrieval success and multi-hop accuracy, especially on complex queries.
Appendix E KG-R1 details
E.1 Training algorithm
@@
Input: Dataset $D=\{(q,G,A_{q})\}$, base LLM $\pi_{0}$, maximum number of turns $H$, rollouts NN
@@
      Global: $R^{\text{global},(n)}=w_{\text{H1}}\cdot\text{H1}(\hat{A}_{q},A_{q})+w_{\text{ret}}\cdot v_{\text{ret}}$;
@@
E.2 Server Instruction Prompt
The server instruction template consists of a set of query actions and descriptions. We use the same server instructions for both training and evaluation across experiments.
@@ E.2 Server Instruction Prompt
E.3 Reward Weights
Our RL objective combines per-turn and episode-level signals with fixed weights. Per-turn rewards encourage well-structured interaction and effective retrieval: $r_{t}^{\mathrm{turn}}=w_{\mathrm{fmt}}\,s_{\mathrm{fmt}}+w_{\mathrm{kg}}\,s_{\mathrm{kg}}+w_{\mathrm{ans}}\,s_{\mathrm{ans}}$, where $s_{\mathrm{fmt}}$ scores output structure/format validity, $s_{\mathrm{kg}}$ rewards schema-valid, non-empty KG queries, and $s_{\mathrm{ans}}$ checks final-answer formatting/consistency. Episode-level reward emphasizes answer correctness and retrieval coverage: $R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}(\hat{A}_{q},A_{q})+w_{\mathrm{ret}}\cdot v_{\mathrm{ret}}$, with $v_{\mathrm{ret}}\in\{0,1\}$ indicating adequate retrieval coverage. Unless otherwise noted, we use the default values listed in Table 6.
@@ E.4 RL Implementation and Hyperparameters
Batch Configuration and Gradient Accumulation. The 3B model uses a training batch size of $128$ with a validation batch size of $256$, whereas the 7B model uses a training batch size of $256$ with a validation batch size of $128$. During GRPO rollout, we collect $N=16$ trajectories per prompt for the 3B model and $N=8$ for the 7B model to balance exploration with memory constraints. The mini-batch size is $128$ for both models, with dynamic batch sizing enabled to utilize GPU memory efficiently.
@@
E.5 KG-R1 Retreival Server Details
Setup. We implemented the KG-R1 retrieval server with FastAPI [21] and Uvicorn [10] to support (i) a schema-free 1-hop KG query API, (ii) high-throughput async batch execution, and (iii) robust validation and observability (regex action parsing, standardized <information> wrapping with auto-closure, and final-turn safeguards). After the KG-R1 agent generates a response, the parsed <kg-query> action is sent to the server, which performs exact string matching over the per-question knowledge graph to resolve entities and relations and returns the retrieved information. If the call is invalid—one of: KG_SERVER_ERROR: Invalid Action; KG_FORMAT_ERROR: Missing Required Fields; KG_FORMAT_ERROR: Wrong Argument Count; KG_SAMPLE_NOT_FOUND: Sample Missing; KG_ENTITY_NOT_FOUND: Entity Not in KG; KG_RELATION_NOT_FOUND: Invalid Relation; KG_NO_RESULTS: No Relations Found; KG_NO_RESULTS: No Entities Found—the server responds with a descriptive error message (see box below).
KG-R1 Server Error Examples KG_SERVER_ERROR: Invalid Action <error>Action "get_entity_info" not available (use: get_head_relations, get_tail_relations, get_head_entities, get_tail_entities)</error>  KG_FORMAT_ERROR: Missing Required Fields <error>Missing required fields for get_tail_entities: relation_name</error>  KG_FORMAT_ERROR: Wrong Argument Count <error>get_tail_relations accepts only one entity argument</error>  KG_SAMPLE_NOT_FOUND: Sample Missing <error>Sample "sample_12345" not found in KG</error>  KG_ENTITY_NOT_FOUND: Entity Not in KG <error>Entity "Barack Obamaa" not found in KG</error>  KG_RELATION_NOT_FOUND: Invalid Relation <error>Relation "location.capital" not found in KG</error>  KG_NO_RESULTS: No Relations Found <error>No tail relations found for entity "Random_Entity_123" in knowledge graph</error>  KG_NO_RESULTS: No Entities Found <error>No tail entities found for relation "film.director.film" with head "Barack Obama" in knowledge graph</error>
@@ E.5 KG-R1 Retreival Server Details
Appendix F Supplementary Experimental Results
F.1 Latency and Throughput Analysis
@@
Batched Throughput We evaluate batched inference at batch size $64$ on a single NVIDIA H100 (Table 7). LLM-only baselines (no KG calls) achieve high throughput—$81.8$ (IO) and $70.1$ (CoT) samples/s—driven by short generation (43.0/206.0 tokens per sample). KG-R1 incurs KG-server retrieval, reducing throughput but remaining practical for offline processing: the single-run setting reaches $3.7$ samples/s ($1205.9$ gen tokens/s) with $4.4$ KG calls per query, while the N=4N{=}4 runs setting trades throughput for more KG interaction ($17.5$ calls per query), yielding $2.0$ samples/s ($612.1$ gen tokens/s). Overall, KG-R1 sustains batch processing on one H100 while supporting KG-grounded reasoning.
@@
Configuration Samples Batch KG Calls Calls/Sample Total (s) Gen Tok./Sample Samples/s Gen Tok./s IO 256 64 0 0.0 12.4 43.0 81.8 887.7 CoT 256 64 0 0.0 14.1 206.0 70.1 3740.1 KG-R1 (single run) 256 64 1127 4.4 73.2 345.0 3.7 1205.9 KG-R1 (N=4N{=}4 runs) 256 64 4478 17.5 142.2 340.0 2.0 612.1
F.2 Ablation Studies
All ablations in Table 3 evaluate KG-R1 with Qwen-2.5-3B-it trained on WebQSP and CWQ using a maximum turn budget of H=5H{=}5. We report the full ablation table in Table 3 and training curves in Figures 7.
Turn Reward. We vary the turn-level reward by setting the weights to $w_{\text{fmt}}{=}0$, $w_{\text{kg}}{=}0$, and $w_{\text{ans}}{=}0$ (the default is $0.5$ for all weights).
Turn-level Advantage. Instead of computing the turn-level group advantage $A_{t}^{(n)}$ in Sec. 3.3, we compute a trajectory-wise group advantage